# Red Team Diffuser: Exposing Toxic Continuation Vulnerabilities in Vision-Language Models via Reinforcement Learning

### v2

@@ 5.1. Main Results
Table 3. Cross-model transfer attack results on Gemini-1.5-flash. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
@@
RTD w/o alignment 4.94 65.29 2.68 38.89 4.19 80.13 80.30
RTD w/ alignment (L) 5.36 64.40 3.09 39.69 4.69 79.06 79.14
RTD w/ alignment (M) 5.27 65.77 3.17 39.74 4.43 81.25 81.33
RTD w/ alignment (H) 5.54 66.63 3.52 39.91 4.45 80.84 81.00
@@
RTD w/o alignment 7.20 63.56 4.43 22.02 6.86 77.97 81.49
RTD w/ alignment (L) 7.19 63.01 4.68 21.50 6.35 76.98 80.83
RTD w/ alignment (M) 7.02 64.29 4.68 22.57 7.52 79.01 82.52
RTD w/ alignment (H) 7.38 64.73 4.86 22.16 7.13 79.09 82.11
Hold-out Set Table 2 presents the results of cross-prompt transfer attacks on the hold-out set against LLaVA. Since these prompts were not seen during training, this evaluation provides a robust measure of the generalization ability of our proposed methods. Interestingly, the hold-out set exhibits a higher average baseline toxicity compared to the original attack set. This is likely due to the inherent nature of the hold-out prompts, which tend to be more inherently toxic and thus more prone to eliciting toxic responses from the target VLM. The Greedy-Prompt Attack demonstrates a clear improvement over the Single-Prompt Attack according to both the Detoxify classifier and the Perspective API. For instance, using Detoxify, the Greedy-Prompt Attack achieved an ‘Any*’ rate of 77.57%, compared to 76.64% for the Single-Prompt Attack. Similarly, with the Perspective API, the Greedy-Prompt Attack achieved an ‘Any*’ rate of 77.60%, compared to 75.63% for the Single-Prompt Attack. More notably, the full Red Team Diffuser (RTD) consistently demonstrates superior performance across most metrics, achieving a peak ‘Any*’ rate of 79.45% according to Detoxify with RTD w/o alignment, as well as notable increases in specific attributes such as Obscene (67.44%) and Insult (42.02%). Furthermore, results from the Perspective API show that RTD w/o alignment achieves the highest ‘Any*’ rate of 79.28%, along with leading scores in attributes such as Profanity (68.13%) and Severe Toxicity (7.12%). These consistent gains clearly validate the strong generalization capability of RTD across diverse, unseen prompts and across different toxicity classifiers, underscoring its effectiveness as a black-box multimodal jailbreak method.
Table 4. Cross-model transfer attack results on Llama-3.2-11B-Vision-Instruct. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
@@
RTD w/o alignment 1.58 32.49 0.16 16.54 2.84 39.93 41.01
RTD w/ alignment (L) 1.67 32.83 0.33 15.37 2.50 41.43 41.85
RTD w/ alignment (M) 1.58 32.63 0.33 16.77 2.25 41.23 41.81
RTD w/ alignment (H) 2.17 31.30 0.16 16.27 2.25 40.56 41.15
@@
RTD w/o alignment 2.67 35.47 1.16 7.26 2.83 39.23 43.90
RTD w/ alignment (L) 2.58 35.30 1.00 8.01 2.58 40.98 45.24
RTD w/ alignment (M) 3.25 34.27 1.41 7.92 3.16 39.69 44.37
RTD w/ alignment (H) 3.58 34.19 0.58 8.34 2.83 39.94 43.78
5.2. Cross-Model Transferability
We evaluate the cross-model transferability of our proposed method by using the RTD trained solely on LLaVA (Liu et al., 2024a). Transfer attacks are performed on two unseen VLMs: Gemini-1.5-flash (Team et al., 2023) and Llama-3.2-11B-Vision-Instruct. Tables 3 and 4 present the toxicity rates evaluated by both the Detoxify classifier and the Perspective API. These results demonstrate that RTD generalizes effectively across models, highlighting its robustness and potential threat as a transferable multimodal jailbreak method.
Table 3 presents the results on Gemini-1.5-flash. Despite being a commercial VLM, Gemini exhibits high baseline toxicity under the Text-only Attack, with ‘Any*’ rates reaching 76.23% (Detoxify) and 78.26% (Perspective), indicating that harmful completions are already prevalent even without adversarial visual input. Furthermore, all RTD variants elevate toxicity levels, with RTD w/ alignment (M) achieving up to 81.33% and 82.52% on Detoxify and Perspective API, respectively. These findings highlight that even commercial VLMs are vulnerable to context-driven multimodal attacks, especially when guided by semantically adversarial images. They also underscore the importance of evaluating alignment in continuation-based and visually conditioned scenarios—not just instruction-following tasks.
@@ 5.2. Cross-Model Transferability
Table 4 reports the transfer results on LLaMA-3.2-11B-Vision-Instruct, a model with more advanced alignment safeguards. Here, the Text-only Attack yields substantially lower toxicity rates—15.02% under Detoxify and 18.68% under Perspective API—reflecting a high refusal rate for toxic continuations. However, RTD is still able to bypass these safeguards. RTD w/ alignment (L) achieves 41.85% (Detoxify) and 45.24% (Perspective), with RTD (M) performing similarly. These results demonstrate RTD’s ability to bypass even stronger alignment mechanisms, transferring its attack effectiveness across models. Together, they highlight the broader risk of multimodal toxicity attacks against both commercial and open-source VLMs, and underscore the need for more robust, context-aware alignment strategies in future VLM design.
@@
Figure 3. Comparison of images generated by the general-purpose Stable Diffusion (left) and the Red Team Diffuser (right). Given the same image prompts, RTD produces images that subtly elicit discomfort or unease, making them more likely to trigger harmful continuations in VLMs.
Table 5. CLIPScore comparison across different diffusers.
Stable Diffusion RTD (w/o) RTD (L) RTD (M) RTD (H)
@@
5.3. Visualizations
Figure 3 presents a qualitative comparison between images generated by the general-purpose Stable Diffusion model (left) and our Red Team Diffuser (RTD) (right). For RTD-generated images, the alignment reward weights $\lambda$ progressively increase from left to right, starting with a weight of 0 in the leftmost column and followed by 0.05, 0.1, and 0.2. Although all models receive the same image prompts, the resulting images diverge markedly in tone and visual implications. While the general-purpose Stable Diffusion consistently produces neutral and benign visuals. In contrast, RTD generates images that subtly evoke discomfort or toxicity-relevant cues. For instance, in the second row, RTD produces cartoonish and exaggerated caricatures that distort the features of an overweight person eating a sandwich, in contrast to the more neutral depiction by Stable Diffusion. In the third row, where the prompt involves generating an image of a woman, RTD creates visuals that are noticeably more sexually suggestive and explicit. Similarly, in the fourth row, RTD introduces additional elements, such as garbage and ominous shadows, into the school hallway scene generated by Stable Diffusion, thereby amplifying the potential for toxic continuations. The observed differences in visual elements and stylistic features underscore RTD’s ability to adapt visual semantics in ways that exploit the vulnerabilities of target VLMs. The comparison highlights how reinforced optimization—not merely prompt selection—enables subtle but impactful visual manipulation aligned with toxic continuation goals.
To complement these qualitative observations, we report CLIPScore (Hessel et al., 2021) as a quantitative measure of image-text semantic alignment. As shown in Table 5, all RTD variants achieve CLIPScores comparable to or higher than Stable Diffusion, indicating that the adversarial images remain semantically faithful to their prompts despite conveying more toxic or discomforting signals. Notably, RTD with the highest alignment reward ($\lambda=0.2$) attains a CLIPScore of 44.80, surpassing the baseline (44.47) and demonstrating that our stealth-aware reward mechanism enhances both visual regularity and the adversarial effectiveness of the generated images.
6. Limitations
@@
While the Red Team Diffuser (RTD) framework shows strong performance and cross-model transferability, it has several limitations: 1. Task specificity. RTD is designed for toxic continuation tasks, and its effectiveness on other multimodal attacks, such as instruction-level jailbreaks, is yet to be explored. 2. Computational cost. The RL-based fine-tuning requires significant GPU resources, which may be impractical for resource-constrained users.
7. Conclusion
We present Red Team Diffuser (RTD), a novel framework for uncovering critical vulnerabilities in Vision-Language Models (VLMs) through semantically adversarial image generation. We formalize cross-modal toxicity amplification as a new attack surface, where visually coherent yet maliciously tuned images exacerbate harmful text continuations—an aspect largely overlooked by current alignment efforts. RTD combines greedy prompt search with reinforcement fine-tuning to generate images that reliably trigger toxic outputs. Experimental results validate its effectiveness: RTD increases LLaVA’s toxicity rate by 10.69% on the original attack set and 8.91% on a hold-out set, while demonstrating strong cross-model transferability, achieving toxicity gains of 5.1% on Gemini and 26.83% on LLaMA-3.2-11B-Vision-Instruct. Qualitative analysis further illustrates how RTD manipulates visual cues to subtly induce toxic continuations. These findings expose substantial blind spots in existing VLM alignment strategies and underscore the need for more robust, multimodal-aware safety mechanisms. In future work, we aim to expand RTD to broader adversarial objectives and model families, offering deeper insights into VLM vulnerabilities and guiding the development of stronger alignment and defense strategies.

### v4

@@ V-B Cross-Model Transferability
TABLE III: Cross-model transfer attack results on Gemini-1.5-flash. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
@@
RedDiff w/o alignment 4.94 65.29 2.68 38.89 4.19 80.13 80.30
RedDiff w/ alignment (L) 5.36 64.40 3.09 39.69 4.69 79.06 79.14
RedDiff w/ alignment (M) 5.27 65.77 3.17 39.74 4.43 81.25 81.33
RedDiff w/ alignment (H) 5.54 66.63 3.52 39.91 4.45 80.84 81.00
@@
RedDiff w/o alignment 7.20 63.56 4.43 22.02 6.86 77.97 81.49
RedDiff w/ alignment (L) 7.19 63.01 4.68 21.50 6.35 76.98 80.83
RedDiff w/ alignment (M) 7.02 64.29 4.68 22.57 7.52 79.01 82.52
RedDiff w/ alignment (H) 7.38 64.73 4.86 22.16 7.13 79.09 82.11
Table III presents the results on Gemini-1.5-flash. Despite being a commercial VLM, Gemini exhibits high baseline toxicity under the Text-Only Attack, with ‘Any*’ rates reaching 76.23% (Detoxify) and 78.26% (Perspective API), indicating that harmful completions are already prevalent even without adversarial visual input. Furthermore, all RedDiffuser variants elevate toxicity levels, with RedDiff w/ alignment (M) achieving up to 81.33% and 82.52% on Detoxify and Perspective API, respectively. These findings highlight that even commercial VLMs are vulnerable to context-driven multimodal attacks, especially when guided by semantically adversarial images. They also underscore the importance of evaluating alignment in continuation-based and visually conditioned scenarios, not just instruction-following attacks.
TABLE IV: Cross-model transfer attack results on Llama-3.2-11B-Vision-Instruct. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
@@
RedDiff w/o alignment 1.58 32.49 0.16 16.54 2.84 39.93 41.01
RedDiff w/ alignment (L) 1.67 32.83 0.33 15.37 2.50 41.43 41.85
RedDiff w/ alignment (M) 1.58 32.63 0.33 16.77 2.25 41.23 41.81
RedDiff w/ alignment (H) 2.17 31.30 0.16 16.27 2.25 40.56 41.15
@@
RedDiff w/o alignment 2.67 35.47 1.16 7.26 2.83 39.23 43.90
RedDiff w/ alignment (L) 2.58 35.30 1.00 8.01 2.58 40.98 45.24
RedDiff w/ alignment (M) 3.25 34.27 1.41 7.92 3.16 39.69 44.37
RedDiff w/ alignment (H) 3.58 34.19 0.58 8.34 2.83 39.94 43.78
@@
Table IV reports the transfer results on LLaMA-3.2-11B-Vision-Instruct, a model with more advanced alignment safeguards [wang2024ideator]. In this setting, the Text-Only Attack produces substantially lower toxicity rates, with 15.02% under Detoxify and 18.68% under the Perspective API, indicating a high refusal rate for toxic continuations. However, RedDiffuser is still able to bypass these safeguards. RedDiff w/ alignment (L) achieves 41.85% (Detoxify) and 45.24% (Perspective API), with RedDiff (M) performing similarly. These results demonstrate RedDiffuser’s ability to bypass even stronger alignment mechanisms, transferring its attack effectiveness across models. Together, they highlight the broader risk of multimodal toxicity attacks against both commercial and open-source VLMs, and underscore the need for more robust, context-aware alignment strategies in future VLM design.
V-C Visualizations
Figure 3 qualitatively compares images generated by Stable Diffusion (left) and RedDiffuser (right). For RedDiffuser, the alignment reward weight $\lambda$ increases from left to right: 0, 0.05, 0.1, and 0.2. Although all models receive the same image prompts, the resulting images diverge markedly in tone and visual implications. The general-purpose Stable Diffusion consistently produces neutral and benign visuals. In contrast, RedDiffuser generates images that subtly evoke discomfort or toxicity-relevant cues. For instance, in the second row, RedDiffuser produces cartoonish and exaggerated caricatures that distort the features of an overweight person eating a sandwich, in contrast to the more neutral depiction by Stable Diffusion. In the third row, where the prompt involves generating an image of a woman, RedDiffuser creates visuals that are noticeably more sexually suggestive and explicit. Similarly, in the fourth row, RedDiffuser introduces additional elements, such as garbage and ominous shadows, into the school hallway scene generated by Stable Diffusion, thereby amplifying the potential for toxic continuations. The observed differences in visual elements and stylistic features underscore RedDiffuser’s ability to adapt visual semantics in ways that exploit the vulnerabilities of target VLMs. The comparison highlights how reinforced optimization, instead of mere prompt selection, enables subtle but impactful visual manipulation aligned with toxic continuation goals. To quantify semantic alignment, we report CLIPScore [hessel2021clipscore] in Table V. All RedDiffuser variants achieve scores on par with or exceeding the baseline (44.47), with RedDiffuser at $\lambda=0.2$ reaching 44.80, demonstrating that our stealth-aware reward maintains prompt relevance while enhancing adversarial potency.
@@ V-C Visualizations
Figure 3: Comparison of images generated by the general-purpose Stable Diffusion (left) and the RedDiffuser (right). Given the same image prompts, RedDiffuser produces images that subtly elicit discomfort or unease, making them more likely to trigger harmful continuations in VLMs.
TABLE V: CLIPScore comparison across different diffusers.
Stable Diffusion RedDiff (w/o) RedDiff (L) RedDiff (M) RedDiff (H)
@@
V-D Extended Evaluation Under External Guardrails
Table VI reports the evaluation results when LLaVA is paired with external image-level guardrails. In this configuration, generated images are first processed by the Stable Diffusion NSFW filter and a VLM-based semantic NSFW checker; images flagged by either checker will be refused by the model and thus do not produce toxic continuations. Comparisons in this section use the RedDiff w/ alignment (M) setting, which provides a strong balance between toxicity amplification and semantic coherence. In addition to reporting toxicity rates as in previous sections, we also measure the Guardrail Pass Rate (GPR), defined as the fraction of generated images that pass both checkers. GPR indicates how often a generated image remains admissible under external safety filtering.
Interestingly, we observe that RedDiff (Base) achieves a higher GPR than the unmodified Stable Diffusion (54.13% vs. 49.96%), despite not being trained with guardrail feedback. This suggests that RedDiff (Base) does not simply increase toxicity by producing NSFW imagery, but instead shifts generation toward more subtle, suggestive visual cues that are less likely to be filtered. With checker-aware training, RedDiff (Guard) further improves both bypass ability and toxicity. Under Detoxify, the ‘Any*’ toxicity rate increases from 38.28% to 69.39%, with GPR rising to 94.50%; similar trends hold under the Perspective API (‘Any*’: 69.47%). These results show that checker-aware optimization enables the model to consistently produce images that pass external filters while still eliciting harmful continuations.
Figure 4 shows visual examples from the three diffusion variants. Stable Diffusion generates relatively neutral imagery. In comparison, RedDiff (base) produces images with stronger and clearer semantic cues—for instance, exaggerated muscular bodies or heightened conflict in the scene. RedDiff (guard) further adjusts these cues to remain accepted under external guardrails. We observe a tendency toward stylized or cartoon-like rendering, softer shapes, and implied rather than directly exposed body features. Such images retain the suggestive cues needed to influence the VLM, while being less likely to trigger NSFW checkers. Overall, these qualitative patterns align with the quantitative results: checker-aware optimization encourages the model to adopt visual styles that are subtle enough to pass safety filtering while still affecting the model’s continuation.
TABLE VI: Black-box attack results on the full dataset against LLaVA with external guardrails. Toxicity rates (%) are evaluated using Detoxify and Perspective API. Any* indicates toxicity in at least one category. Guardrail Pass Rate (GPR, %) measures the fraction of generated images that pass both external safety checkers.
Model Detoxify (%) GPR (%)
Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
Stable Diffusion 3.09 32.28 1.42 18.93 1.75 37.45 37.45 49.96
RedDiff (Base) 3.17 32.19 1.25 18.93 1.50 38.12 38.28 54.13
RedDiff (Guard) 5.67 58.89 2.67 35.03 2.59 69.22 69.39 94.50
Model Perspective API (%) GPR (%)
Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
Stable Diffusion 3.92 32.78 2.17 12.09 2.67 37.11 38.20 49.96
RedDiff (Base) 4.34 33.03 2.50 13.76 2.75 37.78 39.78 54.13
RedDiff (Guard) 7.26 59.47 4.67 20.85 4.42 67.31 69.47 94.50
Figure 4: Visual comparison across diffusion variants. Stable Diffusion outputs relatively neutral imagery; RedDiff (Base) produces stronger semantic cues that influence continuations. RedDiff (Guard) adjusts visual subtlety to consistently pass external safety checkers while retaining adversarial intent.
VI Limitations
@@
While the RedDiffuser framework shows strong performance and cross-model transferability, it has several limitations: 1. Task specificity. RedDiffuser is designed for toxic continuation tasks, and its effectiveness on other multimodal attacks, such as instruction-level jailbreaks, is yet to be explored. 2. Computational cost. The RL-based fine-tuning requires considerable GPU resources, which may pose challenges for users with limited budgets.
VII Conclusion
We introduced RedDiffuser, a novel framework that exposes a previously underexamined vulnerability in Vision-Language Models (VLMs): cross-modal toxic continuation, where visually plausible images can amplify harmful text completions. RedDiffuser combines greedy prompt search with reinforcement learning-based diffusion tuning to generate semantically adversarial images that reliably induce toxic continuations while preserving visual coherence. Our experiments show that RedDiffuser substantially increases toxicity rates on LLaVA (+10.69% on the original set and +8.91% on a hold-out set) and transfers effectively across models, raising toxicity by +5.1% on Gemini-1.5-flash and +26.83% on LLaMA-Vision-Instruct. To reflect more realistic deployment scenarios, we further evaluate attacks under external NSFW guardrails. A checker-aware variant of RedDiffuser significantly improves the guardrail pass rate (GPR), reaching 94.50%, while maintaining high toxicity. This result shows that toxic continuation remains feasible even under image-level safety filtering. These findings reveal fundamental blind spots in current multimodal safety alignment and highlight the necessity of developing defenses that jointly consider visual semantics and continuation behavior. Future work will extend RedDiffuser to broader adversarial objectives and more diverse model architectures to deepen the understanding of multimodal safety risks.